# Optimizing Instructions and Demonstrations for Multi-Stage Language Model Programs

### v1

@@ Abstract
Language Model Programs, i.e. sophisticated pipelines of modular language model (LM) calls, are increasingly advancing NLP tasks, but they require crafting prompts that are jointly effective for all modules. We study prompt optimization for LM programs, i.e. how to update these prompts to maximize a downstream metric without access to module-level labels or gradients. To make this tractable, we factorize our problem into optimizing the free-form instructions and few-shot demonstrations of every module and introduce several strategies to craft task-grounded instructions and navigate credit assignment across modules. Our strategies include (i) program- and data-aware techniques for proposing effective instructions, (ii) a stochastic mini-batch evaluation function for learning a surrogate model of our objective, and (iii) a meta-optimization procedure in which we refine how LMs construct proposals over time. Using these insights we develop MIPRO, a novel optimizer that outperforms baselines on five of six diverse LM programs using a best-in-class open-source model (Llama3-8B), by as much as 13% accuracy. We will release our new optimizers and benchmark in DSPy at http://dspy.ai.
@@ 1 Introduction
Solving complex tasks with Language Models (LMs) often requires applying sophisticated prompting techniques (Wei et al., 2022; Chen et al., 2022) and chaining them together into multi-stage pipelines (Wu et al., 2022; Dohan et al., 2022; Khattab et al., 2022; Yao et al., 2023; Schlag et al., 2023). Such Language Model (LM) Programs continue to advance NLP tasks (Pourreza and Rafiei, 2023; Khattab et al., 2024; Ridnik et al., 2024) through systematic composition (Khattab et al., 2021; Creswell and Shanahan, 2022; Pan et al., 2024) and tool use (Qin et al., 2023). However, LM programs today are commonly designed via “prompt engineering”: crafting lengthy prompts via manual trial and error to coerce a specific LM to operate each step in a specific pipeline.
Recent work such as APE (Zhou et al., 2023), OPRO (Yang et al., 2024), and EvoPrompt (Guo et al., 2024) presents prompt optimizers, i.e. algorithms that search over strings to identify high-performing prompts. Unfortunately, the majority of this work does not directly apply to multi-stage LM programs in which we lack gold labels or evaluation metrics for the individual LM calls. Khattab et al. (2024) study how to express arbitrary LM pipelines such that the instructions, demonstrations (input/output examples), and LM weights of each LM call are treated as parameters that can be algorithmically optimized toward any metric. While the authors present optimizers that can create demonstrations of multi-stage pipelines and use them to optimize prompts and weights, their proposed optimizers cannot tune free-form instructions for sophisticated multi-prompt pipelines.
We seek to efficiently optimize prompts in arbitrary LM programs, especially those with multiple stages (Figure 1) and explore approaches that hold under weak assumptions, consistent with the abstractions from the DSPy programming model (Khattab et al., 2024). In particular, we assume no access to LM weights, log-probabilites, or handwritten metrics or labels for intermediate stages in a chain of LM calls. We require only the LM program itself, a metric to optimize, and a training set of inputs (and, depending on the metric, final outputs). This enables us to optimize LM programs that make use of blackbox LMs without collecting detailed training data for each module.
@@
We make three contributions. First, we present a formalization of the problem of optimizing language model programs (§2) and propose an algorithm design space with three strategies to address the challenge of prompt proposal and three strategies to resolve the issue of credit assignment (§3). Second, we release a benchmark suite for LM program optimizers spanning six tasks (§5). Third, we construct and evaluate a rich subset of possible algorithms for prompt optimization (§4). Highlighted amongst these algorithms is MIPRO (Multi-prompt Instruction PRoposal Optimizer) which outperforms all baseline optimizers on five of six tasks in our benchmark, by as much as 13% accuracy improvement. Using our algorithms, we derive five key lessons for practitioners looking to optimize LM programs (§6).
@@ 3 Designing LM Program Optimizers
Algorithm 1 defines a general optimization framework for LM programs. We seek efficient instantiations of this algorithm in which we minimize the number of times the program is invoked (Line 10) and, as a result, the number of times we must sample proposals (Line 8). This requires efficient search over the use of natural language to directionally influence the behavior of LMs within larger programs. To this end, we are especially interested in building multi-stage LM programs that encode useful strategies for handling the proposal and credit assignment challenges (Section 1). In this section, we present several novel or improved strategies for this. Section 4 then defines a few effective compositions of these strategies that allow us to empirically study their properties in practice.
@@ 3.1 The Proposal Problem
To make the approximate optimization tractable, we must be able to efficiently sample candidate prompts that are properly aligned with the nature of our task, program, data, and metric. To do this, we could (i) construct instructive example traces (i.e. inputs and outputs for each module) that demonstrate how the LM should conduct the task, (ii) collect and summarize important factors that could shape the construction of high-quality instructions, and/or (iii) meta-optimize how LMs are used to propose high-performing instructions.
Bootstrapping Demonstrations Khattab et al. (2022, 2024) study a simple yet surprisingly effective rejection-sampling strategy for optimizing the prompts of LM programs. In essence, they sample inputs from the training set and run to sample input/output traces for each module in the program. If the output scores are successful, i.e. for some threshold and , they treat all values in the trace as a potential labeled demonstration for the respective module in . Given these potential demonstrations, the optimization problem is reduced to selecting combinations of demonstrations (within and across modules) that serve as effective few-shot examples for prompting, an approach that Khattab et al. (2024) find can often outperform hand-written demonstrations for multi-stage programs.
Grounding When asking LMs to craft instructions, how do we define the task for them in a way that encourages proposals that are grounded in the optimization environment? We hypothesize that effective proposals reflect the task dynamics, which can be inferred from properties of the dataset, the program, and successful traces through the program. Hence, we build a zero-shot LM program for (i) characterizing patterns in the raw dataset , (ii) summarizing the program’s control flow, (iii) bootstrapping program demonstrations, and (iv) collecting per-stage prompts that were previously evaluated with their evaluation scores on the train set. We consider supplying each of these pieces as context to ground the LM proposing our instructions. Details on constructing the dataset and program summaries are included in Appendix C.
Learning To Propose Every strategy has several proposer hyperparameters, e.g. the temperature used for instruction generation and whether to ground the proposer with a data summary. Optimal configurations of these hyperparameters may depend in practice on the task, program, and LM. For example, the dataset summary may be essential to a logical reasoning task but may distract a small proposer LM for highly familiar tasks like factoid question answering. Motivated by this, in learning to propose, we parameterize proposal hyperparameters for bootstrapping and grounding, and learn a Bayesian model over several trials of proposing demonstrations and instructions to find what proposal strategy works for a particular program.
@@ 3.2 Credit Assignment
Proposed assignments may be combined in many configurations. To search this space, we must identify the contribution of specific choices to high- or low-scoring trajectories. We propose and explore three solutions for this credit-assignment problem: greedy, surrogate, and history-based.
Greedy As one technique, we consider proposing and evaluating single-stage changes to the LM program separately. This method limits the mis-attribution of errors to incorrect stages, but is inefficient, as changes to some stages may not change program-level performance until other stages are improved first. Our preliminary experimentation with greedy credit assignment demonstrated that it was no more effective than other approaches but it imposed considerably worse time complexity.
Surrogate For credit assignment, we propose the use of a surrogate model to predict the quality of a given combination of variables from past evaluated combinations. This model then improves the quality of combinations proposed by the algorithm. In practice, we use Optuna’s implementation of the Tree Structured Parzen Optimizer to build a Bayesian model over the quality of parameter combinations for the LM program (Akiba et al., 2019; Bergstra et al., 2011). This multivariate variation of TPE models joint contributions between parameter choices to total program quality Falkner et al. (2018), learning differing credit assignments across regions of the combinatorial space.
History-Based We propose the use of historical samples as a method of credit assignment and to inform our future proposals. Following the strategy proposed in OPRO (Yang et al., 2023), we rely upon an LM to learn to assign credit from a history of evaluated trajectories and their scores by including these in the context to the proposer LM. In Section 4, we cover a few ways in which the LLM can be instructed to conduct this credit assignment process more or less explicitly.
@@ 4.1 Bootstrap Random Search
Khattab et al. (2024) achieve exceptional results by generating and selecting task demonstrations using random search. This approach fits into our general framework and serves as a strong baseline in our experiments. Here, every prompt used by has variables over demonstrations . The hyperparameters include the number of demonstrations to use for each module, and the number of total sets to bootstrap and evaluate.
To Initialize, given a set of pairs , where is an input to the program and contains metadata (e.g. empty or final answers), we run . The full trace of each run provides a value for each module-level demonstration variable. If the output scores highly, i.e. for some threshold and given metadata or final label , we assume the demonstrations to be potentially valid and add them to a global store . To Propose, we sample a subset of demonstrations from (according to the hyperparameters) and use these to create a partial assignment from demonstration variables to demonstrations for every module in . To Update, the partial assignment is added to a global store with its evaluation score on the training set (or a dedicated validation split thereof). To ExtractOptimizedSets, a top-scoring assignment in is extracted.
@@ 4.2 Module-Level OPRO
We seek to extend the OPRO algorithm (Yang et al., 2023) to an arbitrary LM program , e.g. with stages embedded in a larger program. We must account for the absence of direct evaluators (and of any labels) for intermediate stages.
We first consider an approach we refer to as module-level OPRO: for a given program stage indexed by , we assume that changes to assignments for other stages will not impact the stage’s contribution to the task-level score, and implicitly use the proxy score where is implied by past assignments. Optimizers for each stage are provided with histories of module-level trajectories with proxy scores only. In this approach, hyperparameters include stage-level meta-instructions, a seed instruction , a maximum history length, a function for choosing scored trajectories from , and stage-level choices over grounding contexts and LM hyperparameters.
To Initialize, the dataset is used to score . Call this global score . Each is added to the global store with as its proxy score. To Propose, an LM is prompted with (module-level) meta-instructions and histories to generate candidate instructions . These define a partial assignment . To Update, each is added to . The history is updated according to the trajectory choice function and maximum history length. To ExtractOptimizedSets, a top-scoring partial assignment is extracted from .
@@ 4.3 MIPRO
Bayesian optimization is known for its robustness to noise, as it effectively incorporates uncertainty into the optimization process Snoek et al. (2012). We therefore propose evaluating over mini-batches of our training data, rather than the full set with each iteration. This allows us to explore and exploit parameter configurations more efficiently.
To Initialize, MIPRO proposes a complete set of instructions per module using the proposal hyperparameters , and it bootstraps a complete set of task demonstrations per module . All latent variables in the Bayesian Model are initialized with a uniform prior. To Propose, we use the sampling rule from the Tree-structured Parzen Estimator (Bergstra et al., 2011) to propose a partial assignment of instructions and demos .
To Update, the partial assignments and are scored on a randomly selected mini-batch of B samples, and the scores are used to update the Estimator’s priors over parameter quality. To ExtractOptimizedSets, every steps, the candidate parameterizations of with the highest mean score over trials is evaluated on the full train-set. At the end of the trials, the highest scoring fully evaluated parameterization is returned as the optimal assignment.
@@ 4.4 Other MIPRO variants
| HoVer | Multi-Hop Claim Verify | Multi-Hop Retrieval | 4 | 4 | Retrieval@21 |
@@ 4.5 Other OPRO variants
Program-Level OPRO provides the proposer LM with a history of full, multi-stage trajectories and relies on it to assign credit for task-level scores to the program stages. While Module-level OPRO embeds the assumptions that there is no inter-assignment dependency and that credit assignment across modules is equal, Program-level OPRO makes the assumption that an LM will successfully complete credit-assignment when provided with long trajectory histories. These are all very strong assumptions. In particular, information contained in histories is likely to be lost as history length grows Liu et al. (2023). In our experiments, we opt for using Module-level OPRO because Program-Level OPRO is more complex and did not appear to provide additional performance gains.
@@ 5.1 Benchmark
We develop six tasks (i.e. six groups of dataset, metric, and LM program) to evaluate LM program optimizers. Table 1 presents our tasks, whose full descriptions, splits, and DSPy program psuedocode are presented in Appendices A, B.1 and E, respectively. We use 500 examples for training, 500 for our development, and a test set of 2k examples (or the full test set for tasks with test sets <2k samples).
@@
We also include two classical classification tasks: Iris (Fisher 1936; flower classification given six real-numbered features) and Heart Disease (Detrano et al. 1989; binary classification given 13 categorical and continuous features). Iris can be nearly solved with a simple set of rules (not provided to the LM), so we seek to test if LM program optimizers can automatically teach LMs a Chain-of-Thought behavior to also perform well on such tasks. In contrast, it may be harder to find a small number of crucial patterns in Heart Disease, and we thus test a program that generates three clinical opinions using Chain-of-Thought LM calls and then generates a final judgment accordingly.
@@
| Optimizer | ScoNe | HotPotQA | HoVer | HotPotQA Cond. | Iris | Heart Disease | ||||||||||
| Train | Dev | Test | Train | Dev | Test | Train | Dev | Test | Train | Dev | Test | Train | Test | Train | Test | |
| Instructions only (0-shot) | ||||||||||||||||
| N/A | 57.0 | 56.2 | 69.1 | 35.4 | 31.8 | 36.1 | 30.2 | 30.8 | 25.3 | 13.8 | 10.5 | 6 | 34.7 | 32 | 23.3 | 26.8 |
| Module-Level OPRO G | 70.0 | 67.4 | 76.1 | 36.0 | 31.7 | 36.0 | 30.0 | 30.0 | 25.7 | – | – | – | – | – | – | – |
| Module-Level OPRO | 69.1 | 67.6 | 73.5 | 41.9 | 36.2 | 39.0 | 37.1 | 38.6 | 32.5 | – | – | – | – | – | – | – |
| 0-Shot MIPRO | 66.3 | 65.2 | 71.5 | 40.2 | 34.2 | 36.8 | 37.7 | 38.4 | 33.1 | 22.6 | 20.3 | 14.6 | 56.8 | 56.7 | 26.8 | 25.8 |
| 0-Shot MIPRO++ | 69.0 | 66.9 | 75.7 | 41.5 | 36.2 | 39.3 | 37.1 | 37.3 | 32.6 | – | – | – | – | – | – | – |
| Demonstrations only (Few-shot) | ||||||||||||||||
| Bootstrap RS | 74.9 | 69.6 | 75.4 | 48.6 | 44.0 | 45.8 | 42.0 | 42.0 | 37.2 | 16.4 | 15.0 | 10.4 | 58.9 | 58.7 | 78.4 | 79.2 |
| Bayesian Bootstrap | 75.4 | 67.4 | 77.4 | 49.2 | 44.8 | 46.2 | 44.6 | 44.7 | 37.6 | – | – | – | – | – | – | – |
| Both (Few-shot) | ||||||||||||||||
| MIPRO | 74.6 | 69.8 | 79.4 | 49.0 | 43.9 | 46.4 | 44.7 | 46.7 | 39.0 | 28.4 | 28.1 | 23.3 | 69.1 | 68.7 | 75.2 | 74.2 |

### v2

@@ Abstract
Language Model Programs, i.e. sophisticated pipelines of modular language model (LM) calls, are increasingly advancing NLP tasks. However, building these pipelines requires crafting prompts that are jointly effective for all modules. We study prompt optimization for LM programs, i.e. how to update these prompts to maximize a downstream metric without access to module-level labels or gradients. To make this tractable, we factorize our problem into optimizing the free-form instructions and few-shot demonstrations of every module and introduce several strategies to craft task-grounded instructions and navigate credit assignment across modules. Our strategies include (i) program-and-data-aware techniques for proposing effective instructions, (ii) a stochastic mini-batch evaluation function for learning a surrogate model of our objective, and (iii) a meta-optimization procedure in which we refine how LMs construct proposals over time. Using these insights we develop MIPRO, a novel optimizer that outperforms baselines on five of seven diverse LM programs using a best-in-class open-source model (Llama3-8B), by as much as 13% accuracy. We have released our new optimizers and benchmark in DSPy at http://dspy.ai.
@@ 1 Introduction
Solving complex tasks with Language Models (LMs) often requires applying sophisticated prompting techniques (Wei et al., 2022; Chen et al., 2022) and chaining them together into multi-stage pipelines (Wu et al., 2022; Dohan et al., 2022; Khattab et al., 2022; Beurer-Kellner et al., 2023; Yao et al., 2023; Schlag et al., 2023). Such Language Model (LM) Programs continue to advance NLP tasks (Pourreza and Rafiei, 2023; Khattab et al., 2024; Ridnik et al., 2024) through systematic composition (Khattab et al., 2021; Creswell and Shanahan, 2022; Pan et al., 2024) and tool use (Qin et al., 2023). However, LM programs today are commonly designed via “prompt engineering”: crafting prompts via manual trial and error to coerce a specific LM to operate each step in a specific pipeline. Recent work such as APE (Zhou et al., 2023), OPRO (Yang et al., 2024), and EvoPrompt (Guo et al., 2024) presents prompt optimizers, i.e., algorithms that search over strings to identify high-performing prompts. Unfortunately, the majority of this work does not directly apply to multi-stage LM programs in which we lack gold labels or evaluation metrics for the individual LM calls. Khattab et al. (2024) study how to express arbitrary LM pipelines such that the instructions, demonstrations (input/output examples), and LM weights of each LM call are treated as parameters that can be optimized toward any metric. While the authors present optimizers that can create demonstrations of multi-stage pipelines and use them to optimize prompts, weights, or even both together Soylu et al. (2024), their proposed optimizers cannot tune instructions for multi-prompt pipelines.
We seek to efficiently optimize prompts in arbitrary LM programs, especially those with multiple stages (Figure 1) and explore approaches that hold under weak assumptions, consistent with the abstractions from the DSPy programming model (Khattab et al., 2024). In particular, we assume no access to LM weights, log-probabilites, or handwritten metrics or labels for intermediate stages in a chain of LM calls. We require only the LM program itself, a metric to optimize, and a training set of inputs (and, depending on the metric, final outputs).
@@
We make three contributions. First, we present a formalization of the problem of optimizing language model programs (§2) and propose an algorithm design space with three strategies to address the challenge of prompt proposal and three strategies to resolve the issue of credit assignment (§3). Second, we release a benchmark suite for LM program optimizers spanning seven tasks (§5). Third, we construct and evaluate a rich subset of possible algorithms for prompt optimization (§4). Highlighted amongst these algorithms is MIPRO (Multi-prompt Instruction PRoposal Optimizer) which outperforms baseline optimizers on five of seven tasks in our benchmark, by as much as 13% accuracy improvement. Using our algorithms, we derive five key lessons for practitioners looking to optimize LM programs (§6).
@@ 3 Designing LM Program Optimizers
Algorithm 1 defines a general optimization framework for LM programs. We seek efficient instantiations of this algorithm in which we minimize the number of times the program is invoked (Line 10) and, as a result, the number of times we must sample proposals (Line 8). To this end, we are especially interested in building LM program optimizers with strategies that handle the proposal and credit assignment challenges discussed in Section 1. In this section, we present several novel or improved strategies for this. Section 4 then defines a few effective compositions of these strategies that allow us to empirically study their properties in practice. We showcase how these components come together to make LM Program optimizers in Figure 4.
@@ 3.1 The Proposal Problem
To make the approximate optimization tractable, we must be able to efficiently sample candidate prompts that are well suited to the nature of our task, program, data, and metric. To do this, we leverage another LM as a ‘proposer’ LM, and consider (i) bootstrapping few-shot examples that demonstrate how to conduct the task, (ii) collecting and summarizing important factors that could inform the construction of high-quality instructions, and/or (iii) meta-optimizing how the proposer LM is used to create high-performing instructions.
Bootstrapping Demonstrations Khattab et al. (2022, 2024) study a simple yet surprisingly effective rejection-sampling strategy for optimizing the prompts of LM programs. Inputs are sampled from the training set, and run through to generate input/output traces for each module in the program. If the output scores as measured by metric are successful, i.e. for some threshold and , they treat all values in the trace as a potential labeled demonstrations (i.e. valid input/output examples) for the respective module in . Given these potential demonstrations, the optimization problem is reduced to selecting combinations of demonstrations (within and across modules) that serve as effective few-shot examples for prompting. Khattab et al. (2024) find this can often outperform hand-written demonstrations for multi-stage programs.
Grounding How can we guide our proposal LM to craft performant instructions for a given module? We hypothesize that providing the proposal LM with relevant context, such as properties of the data, the program, and examples of successful task completions, will allow it to create instructions better suited to the task. Hence, we build a zero-shot LM program for (i) characterizing patterns in the raw dataset , (ii) summarizing the program’s control flow, (iii) bootstrapping program demonstrations, and (iv) collecting per-stage prompts that were previously evaluated with their evaluation scores on the train set. We consider supplying each of these pieces as context to ‘ground’ the LM proposing our instructions. Details on constructing the dataset and program summaries are included in Appendix C.
Learning To Propose Every proposal strategy has several hyperparameters, e.g. the temperature used for instruction generation and whether to ground the proposer with a data summary, program control flow, etc. Optimal configurations of these hyperparameters may depend in practice on the task, program, and proposer LM. For example, the dataset summary may be essential to a logical reasoning task but may distract a small proposer LM for highly familiar tasks like factoid question answering. Motivated by this, in learning to propose, we parameterize proposal hyperparameters, and learn a Bayesian model over several trials to find what proposal strategy works for a given task, program, and LM setup.
@@ 3.2 Credit Assignment
Proposed assignments may be combined in many configurations. To search this space, we must identify the contribution of specific choices to LM program performance. We propose and explore three solutions for this credit-assignment problem: greedy, surrogate, and history-based.
Greedy As one technique, we consider proposing and evaluating single-stage changes to the LM program separately. This method limits the mis-attribution of errors to incorrect stages, but is inefficient as (i) changes must be applied one at a time, and (ii) changes to some stages may not change program-level performance until other stages are improved first. Our preliminary experimentation with greedy credit assignment demonstrated that it was no more effective than other approaches but it imposed considerably worse time complexity.
Surrogate To achieve more efficient credit assignment, we propose the use of Bayesian learning, which is known for it’s ability to efficiently optimize functions with multiple latent variables. In this setup, a surrogate model learns to predict the quality of different parameter combinations from previous evaluations, allowing us to focus future exploration on the promising regions of the search space. In practice, we use Optuna’s implementation of the Tree Structured Parzen Optimizer to build a Bayesian model over the quality of parameter combinations for the LM program (Akiba et al., 2019; Bergstra et al., 2011). This multivariate variation of TPE models joint contributions between parameter choices, allowing us to jointly optimize our program’s variables Falkner et al. (2018). In short, surrogate-based optimization allows us to optimize efficiently over a discrete set of existing parameter proposals. However, a shortcoming of this is that it only allows for optimization over a fixed set of proposals. Learnings from past evaluations cannot be used to improve the proposals themselves.
History-Based Here we make the assumption that given a history of past evaluations, a sufficiently strong LM could perform credit assignment, removing the need for an explicit surrogate model. This would allow us to perform credit assignment and propose improved instructions simultaneously, as the proposer LM could in theory do both at once. Following the strategy outlined in OPRO (Yang et al., 2023), we rely on the proposer LM to learn assigned credit from a history of evaluated instructions and their scores by including these in the context. The proposer LM then outputs a new instruction based on this information. In Section 4, we cover a few ways in which the LM can be instructed to conduct this credit assignment process more or less explicitly.
@@ 4.1 Bootstrap Random Search
Khattab et al. (2024) achieve strong results by generating and selecting task demonstrations for each module using random search. This approach fits into our general framework and serves as a strong baseline in our experiments. This optimization procedure (highlighted in Figure 2) works as follows:
The hyperparameters include: , the number of demonstrations to use for each module in , and , the number of total sets to bootstrap and evaluate. To Initialize, sets of few-shot examples are bootstrapped using the Bootstrapping Demonstrations procedure described in 3.1. An input-output pair is randomly sampled from , where is an input to the program and contains metadata (e.g. empty or final answers). Then, is run, which generates a full trace of the steps that used for example . If the output scores highly, i.e. for some threshold and given metadata or final label , we assume the trace to be correct, and use the inputs / outputs for each module in the trace as candidate few-shot examples. This process is repeated until sets of few-shot examples for each module have been bootstrapped. To Propose, we sample a set of few-shot examples and use them to parameterize each module in . To Update, the parameterized is added to a global store along with its evaluation score on the training set (or a dedicated validation split thereof). To ExtractOptimizedSets, the top-scoring assignment is returned.
@@ 4.2 Module-Level OPRO
We seek to extend the OPRO algorithm (Yang et al., 2023) to an arbitrary LM program , e.g. with stages embedded in a larger program. In OPRO, the proposer LM is provided with a history of proposed instructions and their scores, allowing it to learn to propose better instructions over time. To extend OPRO, we first consider an approach we refer to as Module-Level OPRO, in which we assume that the program score is a good enough proxy for an individual instruction’s quality. In other words, even though the program is parameterized with instructions, we will assume that the score is reflective enough of each. Module-level OPRO (Figure 3) works as follows:
To Initialize, a seed instruction is used to parameterize each module in , which is evaluated. To Propose, the resulting score and the module’s instruction are inputted into the proposer LM to create a new instruction for module . This is done for all modules. The set of generated instructions are then used to parameterize , and the parameterized program is evaluated. To Update, each instruction and the score is added to the history for each module. The OPRO sub-routine is run again with the updated histories to create a new set of instructions. We note that optimizers for each stage are provided with histories of module-level trajectories and proxy scores only. This process continues until the maximum number of iterations is reached. To ExtractOptimizedSets, the parameterization of that scored highest is returned.
@@ 4.3 MIPRO
Bayesian optimization is known for its robustness to noise, as it effectively incorporates uncertainty into the optimization process Snoek et al. (2012). We therefore propose evaluating over mini-batches of our training data, rather than the full set with each iteration. This allows us to explore and exploit parameter configurations more efficiently. The MIPRO algorithm (Figure 4) is as follows:
To Initialize, MIPRO bootstraps a set of few-shot example sets and instructions per module using the Bootstrap Demonstration and Grounding strategies respectively, found in 3.1. Latent categorical variables representing the choice of few-shot examples and instructions for each module are initialized with a uniform prior. To Propose, we use the sampling rule from the Tree-structured Parzen Estimator (Bergstra et al., 2011) to select the instructions and few-shot examples used to parameterize . To Update, this parameterized is scored on a randomly selected mini-batch of B samples, and the scores are used to update the Estimator’s priors over parameter quality. To ExtractOptimizedSets, every steps, the candidate parameterizations of with the highest mean score over trials is evaluated on the full train-set. At the end, the highest scoring fully evaluated parameterization is returned as the optimal assignment.
@@ 4.4 Other MIPRO variants
| Iris-Typo | Classification | Chain of Thought | 1 | 1 | Accuracy |
@@
| HoVer | Multi-Hop Claim Verify | Multi-Hop Retrieval | 4 | 4 | Recall@21 |
@@ 4.5 Other OPRO variants
Program-Level OPRO provides the proposer LM with a history of full, multi-stage trajectories and relies on it to assign credit for task-level scores to the program stages. While Module-level OPRO embeds the assumptions that there is no inter-assignment dependency and that credit assignment across modules is equal, Program-level OPRO assumes that an LM will successfully complete credit-assignment when provided with long trajectory histories. These are all very strong assumptions. In particular, information contained in histories is likely to be lost as history length grows Liu et al. (2023). In our experiments, we opt for using Module-level OPRO because Program-Level OPRO is more complex and did not appear to provide additional performance gains.
@@ 5.1 Benchmark
We develop seven tasks (i.e. seven groups of dataset, metric, and LM program) to evaluate LM program optimizers. Table 1 presents our tasks, whose full descriptions, splits, and DSPy program pseudocode are presented in Appendices A, B.1 and E, respectively. We use 500 examples for training, 500 for our development, and a test set of 2k examples (or the full test set if smaller).
@@
We also include two classical classification tasks: Iris (Fisher 1936; flower classification given six real-numbered features) and Heart Disease (Detrano et al. 1989; binary classification given 13 categorical and continuous features). We test Iris in two settings, one with a misspelling in the prompt that may confuse the LM and one corrected.111The misspelling was initially accidental (asking to classify a flower as “versicolour” rather than “versicolor”). However, it serves as a realistic test for optimization from a misspecified prompt, so we report on both settings. Future work should permit LM programming abstractions to detect such errors. Iris can be nearly solved with a simple set of rules (not provided to the LM), so we seek to test if LM program optimizers can automatically teach LMs a Chain-of-Thought behavior to also perform well on such tasks. In contrast, it may be harder to find a small number of crucial patterns in Heart Disease, and we thus test a program that generates three clinical opinions using Chain-of-Thought LM calls and then generates a final judgment accordingly.